OpenStudy (anonymous):

Multiply: (x-4)(x+3) i believe the answer is x^2-x +12 so just checking my answer

OpenStudy (luigi0210):

Might wanna check again.

OpenStudy (luigi0210):

Since it's -4 and +3 it should have a - number

OpenStudy (luigi0210):

Other than that everything else checks out okay.

OpenStudy (anonymous):

oh i messed up when i was typing it

OpenStudy (austinl):

FOIL First Outer Inner Last

OpenStudy (austinl):

\[(a + b)(c + d) = ac + ad + bc + bd\]

OpenStudy (anonymous):

Wrong...

OpenStudy (austinl):

\[(x-4)(x+3) = (x\times x) + (3\times x) - (4\times x) + (3\times -4)= x^2 +3x-4x -12\rightarrow x^2 - x -12\]

OpenStudy (austinl):

**\[\rightarrow x^2 -x -12\]